how to convert si units ti cgs units and vice versa

Grade:10

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

Conversion Table cgs / SI- Units

  cgs Conv.- Factor SI
Distance cm =10-2 m
Mass g =10-3 kg
Time sec = sec
Force dyne =10-5 Newton
Energy erg =10-7 Joule
Charge esu =3.336.10-10 Coulomb
El. Potential statvolt =299.8 Volt
Magnetic Field Gauss =10-4 Tesla
  Further units:
1 eV = 1.602.10-12 erg
1 Ry = 13.6 eV (ionization energy of hydrogen)
1 eV = k.11605oK

 

We are all IITians and here to help you in your IIT JEE preparation.

All the best.

 If you like this answer please approve it....

win exciting gifts by answering the questions on Discussion Forum

Sagar Singh

B.Tech IIT Delhi